OpenStudy (lexy148):

A 0.5 kg ball moves in a circle that is 0.4 m in radius at a speed of 4.0 m/s. Calculate its centripetal acceleration. ac=40 m/s2 now how do you find centripital force?

OpenStudy (surry99):

F= mv^2/r

OpenStudy (lexy148):

thx the answer is 20N

OpenStudy (surry99):

yes...

OpenStudy (lexy148):

nd is the ac part correct

OpenStudy (surry99):

yes...ac=v^2/y = (4 m/s)^2/.4m = 40 m/s^2

OpenStudy (lexy148):

ok thx

OpenStudy (surry99):

oops minor typo...ac=v^2/r

OpenStudy (lexy148):

lol i know what you meant